The average annual electricity consumption by a household refrigerator has decreased from 1800 kWh in 1974 to 450 kWh today. Consider a country with 10 million households with a market penetration of 100 percent for modern refrigerators. If the refrigerators in this country were to continue to consume electric power at the 1974 levels, how much of additional installed power would be needed to meet this extra demand? Assume that the load factor (average fraction of installed power load that is actually produced) of the power plants is 0.8 and the power plants operate 90 percent of the time on an annual basis. 


1
Expert's answer
2021-08-04T02:52:01-0400

Assuming that there were 5 million households in 1974

per household 1,800,000 Wh/5 million = 0.36Whr per household

For the 10 million households, 3600kWh (0.36 × 10 million) is used


the total efficiency of producing electricity = 0.8 × 90% = 0.72


100% = ?

72% = 3600kWh


The power needed = 3600/0.72 = 5000kWh

= 5MWh